OpenStudy (anonymous):

if x is in the first and y in the third quadrant, sinx=3/5, and cosy=-5/13, find the exact value of sin(x+y), cos(x+y), and tan(x+y)?

OpenStudy (campbell_st):

well you need a simple quadrant drawing with some additional information  once you have a and b, then you can use the expansions for sin, cos and tan then substitute the necessary exact values... and simplfy

OpenStudy (anonymous):

how to find a & b?

OpenStudy (campbell_st):

do you know pythagoras' theorem..?

OpenStudy (anonymous):

yes I do

OpenStudy (campbell_st):

ok... you can find a and b using pythagoras, and as it says in the drawing, they are both negative, because of where they are on the number plane. then when you have them for x you can now find cos and tan and for y you can find sin and tan then you can get the expansions...

OpenStudy (anonymous):

can u help me more :(

OpenStudy (anonymous):

\[a^2+3^2=5^2 \therefore a=4, \[\therefore b=12\]

OpenStudy (anonymous):

I did that and next what ?

OpenStudy (anonymous):

cos(x+y) = cos(x)cos(y) - sin(x)sin(y)

OpenStudy (anonymous):

?

OpenStudy (anonymous):

and then ?

OpenStudy (anonymous):

just find cos(x),sin(x),sin(y),cos(y) ...... using the formula as sin = opp/hyp, cos=adj/hyp..... finally tan(x+Y)= sin(x+y)/cos(x+y) ..... u wl get the answer ...

OpenStudy (anonymous):

what is the x and what is y ??

OpenStudy (anonymous):

sin(X)=3/5, cos(x)=-4/5, sin(y)=12/13, cos(y)= -5/13

OpenStudy (anonymous):

the diagram seems to be wrong i think ....